OpenStudy (anonymous):

Suppose a college student pays $1200 for tuition fees. However, she also has to pay $400 for her textbooks (ouch!). What percent of her total education costs does she pay for her books? I NEED HELP PLEASE!!

OpenStudy (anonymous):

400 dollars is what i pay for text books

OpenStudy (anonymous):

400/1600

OpenStudy (anonymous):

however total education costs makes me assume that it needs a variable.....

OpenStudy (anonymous):

so if you divide them by 100 you get 4/16 and if you divide by 4 you get 1/4 so 25%

OpenStudy (anonymous):

SO what is the answer?

OpenStudy (anonymous):

Is it 1600 divided by 400 which is .25?

OpenStudy (anonymous):

400/1600

OpenStudy (anonymous):

= .25

jimthompson5910 (jim_thompson5910):

400/(1200+400) = 400/1600 = 0.25, which is 25% So the student is spending 25% of their total education on books